OpenStudy (anonymous):

A homemaker decides to start a small business making earrings and selling them on the internet. originally, he sets the price of the earnings at $16 a pair.Then he decides to increase the price. The equation: y=-1(x^2-26x-672) describes the profit,y, in dollars , based on the increase in price,x, in dollars. A)write the trinomial in factored form b)determine the x intercepts of the equation c)determine the equation of the axis of symmetry

OpenStudy (anonymous):

brooooo ive been trolled

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

whats up whered the answer go?

OpenStudy (anonymous):

I deleted it.

OpenStudy (anonymous):

Vertex form: y = a(x - h)^2 + k, where (h, k) is the vertex and the equation x = h is the axis of symmetry a) y = -(x^2 - 26x - 672) y = -(x - 42)(x + 16) <=== b) 0 = -(x - 42)(x + 16) x - 42 = 0 or x + 16 = 0 x = 42 or x = -16 x-intercepts: (42, 0), (-16, 0) <=== c) y = -(x^2 - 26x - 672) y = -x^2 + 26x + 672 y = -(x^2 - 26x) + 672 y = -(x^2 - 26x + 169) + 672 + 169 y = -(x - 13)^2 + 341 axis of symmetry: x = 13 <=== d) (13, 341) f) The vertex represents the maximum profit that the business can get and the increase in price of the earrings that makes that profit. In this case, increasing the price by $13 makes a maximum profit of $341. This is an answer to one of my previous questions... So yeah.

OpenStudy (anonymous):

ok thanks i got another question gimme a min

OpenStudy (anonymous):

Well, I would love a medal and fan..aha!

OpenStudy (anonymous):

how do i give u a medal

OpenStudy (anonymous):

What's your question?